# The non-linear dynamics of axion inflation: a detailed lattice study

### v1

@@ II Axion inflation
In this section we provide an overview of the axion inflation model. Using (,+,+,+)(-,+,+,+)( - , + , + , + ) metric signature and the reduced Planck mass $m_{p}\simeq 2.435\cdot 10^{18}$ GeV, we consider an action $S_{\rm tot}=S_{\rm g}+S_{\rm m}$, with $S_{\rm g}~{}\equiv~{}\int{\rm d}x^{4}\sqrt{-g}\frac{1}{2}m_{p}^{2}R$ the standard Hilbert-Einstein term for gravity, and a matter action given by
@@
where ˙$\dot{}\equiv\partial/\partial t$ are derivatives with respect to cosmic time $t$, $a(t)$ is the scale factor, and $H=\dot{a}/a$ is the Hubble rate. We define the electric field $\vec{E}$ in the temporal gauge $A_{0}=0$, and the magnetic field $\vec{B}$, as
@@
with delimited-⟨$\langle$..$\rangle$ … ⟩ indicating volume averaging. Labels K, G, V denote the kinetic, gradient and potential energy densities of the inflaton, respectively, while EM denotes the electromagnetic energy density associated to $A_{\mu}$. We note that Eq. (10) represents the Hubble constraint.
@@ II.1 Linear regime: chiral instability
We choose a chiral vector basis $\{\vec{\varepsilon}^{\,+}(\hat{\mathbf{k}})\,,\vec{\varepsilon}^{\,-}(\hat{\mathbf{k}})\}$, which satisfies the properties
$\begin{split}\hat{\mathbf{k}}\cdot\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})=0\,,\quad\hat{\mathbf{k}}\times\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})=-i\lambda\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})\,,\\ \varepsilon_{i}^{\,\lambda}(\hat{\mathbf{k}})^{*}=\varepsilon_{i}^{\,\lambda}(-\hat{\mathbf{k}})\,,\quad\vec{\varepsilon}^{\,\lambda^{\prime}}(\hat{\mathbf{k}})\cdot\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})^{*}=\delta_{\lambda\lambda^{\prime}}\,.\end{split}$ (19)
We promote the Fourier amplitude into a quantum operator $A^{\lambda}(\tau,{\bf k})\rightarrow\hat{A}^{\lambda}(\tau,{\bf k})\equiv\hat{a}_{{\bf k}}\mathcal{A}^{\lambda}(\tau,{\bf k})$ + $\hat{a}_{-{\bf k}}^{\dagger}\mathcal{A}^{\lambda}(\tau,-{\bf k})^{*}$ by means of standard annihilation and creation operators that verify $[\hat{a}_{{\bf k}},\hat{a}_{{\bf k}^{\prime}}^{\dagger}]\equiv\delta^{(3)}({\bf k}-{\bf k}^{\prime}).$ Using the slow-roll condition $\tau\simeq-1/(aH)$, the mode functions $\mathcal{A}^{\pm}(\tau,{\bf k})$ follow the equation
@@
$\vspace*{2mm}{\rm with}$ $\hskip 51.21504pt\xi\equiv\frac{|\dot{\phi}|}{2H\Lambda}\,.$ (21)
For values $\xi\gtrsim\mathcal{O}(1)$, one of the chiral modes develops an exponential growth, while the other remains in vacuum. For $\dot{\phi}<0$ ($>0$), it is $A^{+}$ ($A^{-}$) the one that experiences the instability. Considering $\xi$ to be approximately constant, as expected deep inside the slow-roll dynamics, a general solution for Eq. (20) is given by
@@
with $F_{0}$ and $G_{0}$ the regular and irregular Coulomb wave functions [11, 12]. The coefficients have been chosen so that for scales deep inside the Hubble radius, $-k\tau\gg 1$, both chiralities are in the so-called Bunch-Davies (BD) vacuum,
@@
Without loss of generality, we choose $\dot{\phi}<0$, so that the excited mode is $\mathcal{A}^{+}$. In the limit $-k\tau\ll 2\xi$, the solution (22) for this mode can be approximated by
@@
Depending on the evolution of $\xi$, the exponential growth of the unstable mode may become sufficiently large so that its contribution to the expansion of the universe (in the form of electromagnetic energy density) may become gradually more and more relevant. Furthermore, the term $\,\vec{E}\cdot\vec{B}\,$ in the $rhs$ of Eq. (5) may also become relevant, affecting the inflaton dynamics. In other words, for sufficiently large coupling $\alpha_{\Lambda}$, as inflation proceeds, there is always a point at which the backreaction-less approximation becomes unreliable. We must, therefore, refine the treatment of the dynamics, as we do next, in order to incorporate the backreaction of the gauge field.
@@ II.3 Local backreaction
$H({\pi}_{\phi}^{\prime}+3\pi_{\phi})$ $=$ $\frac{1}{a^{2}}\nabla^{2}\phi-m^{2}\phi+\frac{\alpha_{\Lambda}}{a^{3}m_{p}}\vec{E}\cdot\vec{B}\,,$ (31)
$H\big{(}\vec{E}^{\prime}+\vec{E}\,\big{)}$ $=$ $\frac{1}{a^{2}}\vec{\nabla}\times\vec{B}-\frac{\alpha_{\Lambda}}{am_{p}}\big{(}\pi_{\phi}\vec{B}-\vec{\nabla}\phi\times\vec{E}\,\big{)},$ (32)
$-3HH^{\prime}$ $=$ ${1\over m_{p}^{2}}(3\rho_{\rm K}+\rho_{\rm G}+2\rho_{\rm EM})\,,$ (33)
@@ II.4 Chiral Projection
which, by construction, is antisymmetric, $\Sigma_{ij}=-\Sigma_{ji}$, and has $\vec{\varepsilon}^{\,\pm}$ as eigenvectors, with eigenvalues ±$\pm 1$, i.e. $\Sigma_{ij}\varepsilon_{j}^{\lambda}=\lambda\varepsilon_{i}^{\lambda}$. Extending the basis to a full orthonormal triad $\{\vec{\varepsilon}^{\,+},\vec{\varepsilon}^{\,-},\hat{k}\}$, and relabelling the vectors as $\vec{\varepsilon}^{\,(-1)}\equiv\vec{\varepsilon}^{\,-},\vec{\varepsilon}^{\,(0)}\equiv\hat{k},\vec{\varepsilon}^{\,(+1)}\equiv\vec{\varepsilon}^{\,+}$, we see that $\vec{\varepsilon}^{\,(0)}$ is also an eigenvector of $\Sigma_{ij}$ with vanishing eigenvalue, i.e. $\Sigma_{ij}\varepsilon_{j}^{\,(0)}=0$. We can therefore write
@@
$\Pi^{\pm}_{ij}(\hat{\mathbf{k}})$$\equiv$ $\frac{1}{2}\left[(\Sigma^{2}(\hat{\mathbf{k}}))_{ij}\pm\Sigma_{ij}(\hat{\mathbf{k}})\right]$ (39)
$=$ $\frac{1}{2}\left(P_{ij}(\hat{\mathbf{k}})\pm\Sigma_{ij}(\hat{\mathbf{k}})\right)\,,$
where in the second line we have used $(\Sigma^{2}(\hat{\mathbf{k}}))_{ij}\equiv\Sigma_{il}(\hat{\mathbf{k}})\Sigma_{lj}(\hat{\mathbf{k}})=P_{ij}(\hat{\mathbf{k}})$, with $P_{ij}(\hat{\mathbf{k}})\equiv\delta_{ij}-\hat{k}_{i}\hat{k}_{j}$ the transverse projector.It then follows that $\Pi^{\lambda}_{ij}(\hat{\mathbf{k}})A_{j}^{\parallel}({\mathbf{k}})=0$ and $\Pi^{\lambda}_{ij}(\hat{\mathbf{k}})A_{j}^{\lambda^{\prime}}({\mathbf{k}})={1\over 2}(1+\lambda\lambda^{\prime})A_{i}^{\lambda^{\prime}}({\mathbf{k}})\equiv A_{i}^{\lambda}({\mathbf{k}})\delta_{\lambda\lambda^{\prime}}$. From here we arrive at the desired property for a chiral projector: given the Cartesian components of a vector, say the gauge field $\vec{A}=(A_{1},A_{2},A_{3})$ or the electric field $\vec{E}=(E_{1},E_{2},E_{3})$, we can obtain their chiral components by means of the projection operation
$A^{\pm}_{i}({\mathbf{k}})\equiv\Pi^{\pm}_{ij}(\hat{\mathbf{k}})A_{j}({\mathbf{k}})\,,~{}~{}E^{\pm}_{i}({\mathbf{k}})\equiv\Pi^{\pm}_{ij}(\hat{\mathbf{k}})E_{j}({\mathbf{k}})\,.$ (40)
@@ III Lattice formulation
We consider that $\phi$ lives on the lattice sites n, $A_{i}$’s on the links (i.e. halfway in between lattice sites, at $\textbf{n}+\hat{\imath}/2$), and we define spatial lattice derivatives as usual, using forward/backward operators $\Delta^{\pm}_{i}\varphi\equiv\pm\frac{1}{\delta x}(\varphi_{\pm i}-\varphi)$, where $\delta x$ is the lattice spacing, with the ±$\pm{i}$ subscripts indicating spatial displacements in one lattice unit in the direction $\hat{\imath}$. We define electric and magnetic fields on the lattice as
@@
$\hskip 142.26378pt-\left.\frac{\alpha_{\Lambda}}{4am_{p}}\sum_{\pm}\sum_{j,k}\epsilon_{ijk}\left\{\left[(\Delta_{j}^{\pm}\phi)E^{(2)}_{k,\pm\hat{\jmath}}\right]_{+\hat{\imath}}+\left[(\Delta_{j}^{\pm}\phi)E^{(2)}_{k,\pm\hat{\jmath}}\right\}\right]\right)\,,$
@@
$\begin{gathered}\rho^{\rm L}_{\rm K}\equiv\frac{1}{2}\left\langle\pi^{2}_{\phi}\right\rangle_{\rm V}\;,\quad\rho^{\rm L}_{\rm G}\equiv\frac{1}{2a^{2}}\Big{\langle}(\vec{\Delta}^{+}\phi)^{2}\Big{\rangle}_{\rm V}\;,\\ \rho^{\rm L}_{\rm V}\equiv\frac{1}{2}m^{2}\left\langle\phi^{2}\right\rangle_{\rm V}\;,\quad\rho^{\rm L}_{\rm EM}\equiv\frac{1}{2a^{4}}\Big{\langle}a^{2}\vec{E}^{2}+\vec{B}^{2}\Big{\rangle}_{\rm V}\;,\end{gathered}$ (49)
with $\langle...\rangle_{\rm V}\equiv\frac{1}{N^{3}}\sum_{\mathbf{n}}(...)$ denoting volume averaging over a lattice of $N$ points per dimension. As long as $N$ is sufficiently large to guarantee that – given a lattice spacing $\delta x$ – the (comoving) lattice length $L=N\delta x$ is much bigger than the (comoving) scales $\lambda_{*}\sim 2\pi/k_{*}$ of the excited fields (say $k_{*}$ is the typical peak scale of the fields’ power spectra), this procedure leads to a well-defined notion of a homogeneous and isotropic expanding background, within the given (comoving) volume $V=L^{3}$.
We have used the public package ${\mathcal{C}}$${\mathcal{L}}$attice [93, 94, 95] as the computational environment to implement our lattice formulation.
@@ III.1 The reciprocal lattice
While different choices of lattice operators are possible to represent a given continuum derivative operation, each choice implies a different lattice momentum ${\bf k}_{\rm L}(\tilde{\bf n})$, that will depart differently from the continuum momenta ${\bf k}(\tilde{\bf n})$ in the ultraviolet (UV) scales ($|\tilde{\bf n}|>N/2$) of the lattice. The Fourier transform of any lattice derivative operator over a function $f$ satisfies the relation
$(\nabla_{i}f)(\tilde{\textbf{n}})=-ik_{\text{L},i}(\tilde{\textbf{n}})f(\tilde{\textbf{n}})\;,$ (51)
which defines the lattice momentum $k_{\text{L},i}(\tilde{\textbf{n}})$ ascribed to the lattice representation of $\nabla_{i}$. In our case, we use backward/forward operators which, if interpreted as centred halfway of the lattice sites, at ${\bf n}+\hat{\imath}/2$, lead to the lattice momentum (both for forward and backward operations)
@@
where $\lambda=\pm$ and $k_{\text{L}}=(k_{\text{L},1}^{2}+k_{\text{L},2}^{2}+k_{\text{L},3}^{2})^{1/2}$. Given the Cartesian components of a vector on the lattice, $A_{i}(\tilde{\bf n})$, we will obtain its chiral components $\lambda=\pm$ as
$A_{j}^{\lambda}(\tilde{\textbf{n}})=\Pi^{{\rm L},\lambda}_{ij}(\tilde{\textbf{n}})A_{j}(\tilde{\textbf{n}})\,.$ (54)
@@
$\vec{\varepsilon}_{{\rm L}}^{\,\pm}(\tilde{\bf n})$$\equiv$ $[\vec{u}(\hat{\mathbf{k}}_{\rm L})\pm i\vec{v}(\hat{\mathbf{k}}_{\rm L})]/\sqrt{2}\,,$ (56)
@@
$\langle{f}^{2}\rangle=\int d\log k~{}\Delta_{\tt f}(k)\,,~{}~{}~{}\Delta_{f}(k)=\frac{k^{3}}{2\pi^{2}}\mathcal{P}_{f}(k)\,,$
@@
As explained in [97], there are different approaches to define the power spectrum on a lattice that mimic $\Delta_{f}({\bf k})$. In this work we opt to use ${\mathcal{C}}$${\mathcal{L}}$attice’s Type I-Version 1 power spectra and canonical binning. This means that in a lattice with $N$ sites per dimension and lattice spacing $\delta x$, there are $l_{\rm max}=[\sqrt{3}N/2]$ bins, labelled as $l=1,2,3,...,l_{\rm max}$, with representative momentum of each bin given by $k_{l}=k_{\rm IR}l$, $k_{\rm IR}\equiv 2\pi/(N\delta x)$. The lattice power spectrum is then
@@
with #$\#_{l}$ the exact multiplicity of modes within the $l$-th bin, and $\langle\left|f(\tilde{\bf n}^{\prime})\right|^{2}\rangle_{R(l)}\equiv{1\over\#_{l}}\sum_{\tilde{\bf n}^{\prime}\in R(l)}\left|f(\tilde{\bf n}^{\prime})\right|^{2}$ an angular average over all modes $\tilde{\bf n}^{\prime}$ inside the $l$-th bin, represented by $R(l)\equiv\{\tilde{\bf n}^{\prime}~{}/~{}|\tilde{\bf n}^{\prime}|\in[l-1/2,l+1/2)\}$. We note that within the IR region ($l<N/2$) of the lattice, Eq. (60) is well approximated by $\Delta_{f}(k_{l})\simeq(k_{l}^{3}/2\pi^{2})(\delta x^{3}/N^{3})\large\langle|f(\tilde{\bf n}^{\prime})|^{2}\large\rangle_{R(l)}$, resembling closely the continuum form $(k^{3}/2\pi^{2})\mathcal{P}_{f}(k)$. Furthermore, the expectation value obtained with the lattice power spectrum just defined, $\left\langle f^{2}\right\rangle=\sum_{l=1}^{l_{\rm max}}{1\over l}\Delta_{f}(k_{l})$, is exactly identical to the lattice volume average $\left\langle f^{2}\right\rangle_{V}\equiv\frac{1}{N^{3}}\sum_{\bf n}f^{2}({\bf n})$, and mimics the continuum counterpart $\int d\log k~{}\Delta_{f}(k)$ as desired, since ${1/l}\equiv{k_{\rm IR}/k_{l}}\simeq dk/k\equiv d\log k$. In practice, we compute power spectra in our simulations for $\phi,\vec{A}^{\pm},\vec{A},\vec{E}^{\pm},\vec{E},\vec{B}$ as $\Delta_{\phi},\Delta_{A}^{(\pm)},\Delta_{A},\Delta_{E}^{(\pm)},\Delta_{E},\Delta_{B}$ via Eq. (60) with $|f(\tilde{\bf n})|^{2}=$ $|\phi(\tilde{\bf n})|^{2}$, $|\vec{A}^{\pm}(\tilde{\bf n})|^{2}$, $|\vec{A}(\tilde{\bf n})|^{2}$, $|\vec{E}^{\pm}(\tilde{\bf n})|^{2}$, $|\vec{E}(\tilde{\bf n})|^{2}$, and $|\vec{B}(\tilde{\bf n})|^{2}$, respectively, with the electric and magnetic fields given by Eq. (42).
For a discussion on other choices of power spectra representations on a lattice, we point the reader to ${\mathcal{C}}$${\mathcal{L}}$attice’s technical note I [97].

### v2

@@ II Axion inflation
In this section we provide an overview of the axion inflation model. Using (,+,+,+)(-,+,+,+) metric signature and the reduced Planck mass $m_{p}\simeq 2.435\cdot 10^{18}$ GeV, we consider an action $S_{\rm tot}=S_{\rm g}+S_{\rm m}$, with $S_{\rm g}~\equiv~\int{\rm d}x^{4}\sqrt{-g}\frac{1}{2}m_{p}^{2}R$ the standard Hilbert-Einstein term for gravity, and a matter action given by
@@
where $\dot{}\equiv\partial/\partial t$ are derivatives with respect to cosmic time $t$, $a(t)$ is the scale factor, and $H=\dot{a}/a$ is the Hubble rate. We define the electric field $\vec{E}$ in the temporal gauge $A_{0}=0$, and the magnetic field $\vec{B}$, as
@@
with $\langle$$\rangle$ indicating volume averaging. Labels K, G, V denote the kinetic, gradient and potential energy densities of the inflaton, respectively, while EM denotes the electromagnetic energy density associated to $A_{\mu}$. We note that Eq. (10) represents the Hubble constraint.
@@ II.1 Linear regime: chiral instability
[…]
$\begin{split}\hat{\mathbf{k}}\cdot\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})=0\,,\quad\hat{\mathbf{k}}\times\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})=-i\lambda\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})\,,\\ \varepsilon_{i}^{\,\lambda}(\hat{\mathbf{k}})^{*}=\varepsilon_{i}^{\,\lambda}(-\hat{\mathbf{k}})\,,\quad\vec{\varepsilon}^{\,\lambda^{\prime}}\hskip-2.84526pt(\hat{\mathbf{k}})\cdot\vec{\varepsilon}^{\,\lambda}(\hat{\mathbf{k}})^{*}=\delta_{\lambda\lambda^{\prime}}\,.\end{split}$ (19)
We promote the Fourier amplitude into a quantum operator $A^{\lambda}(\tau,{\bf k})\rightarrow\hat{A}^{\lambda}(\tau,{\bf k})\equiv\hat{a}_{{\bf k}}\mathcal{A}^{\lambda}(\tau,{\bf k})$ + $\hat{a}_{-{\bf k}}^{\dagger}\mathcal{A}^{\lambda}(\tau,-{\bf k})^{*}$ by means of standard annihilation and creation operators that verify $\hat{a}_{{\bf k}},\hat{a}_{{\bf k}^{\prime}}^{\dagger}]\equiv\delta^{(3)}({\bf k}-{\bf k}^{\prime}).$ Using the slow-roll condition $\tau\simeq-1/(aH)$, the mode functions $\mathcal{A}^{\pm}(\tau,{\bf k})$ follow the equation
@@
$\vskip 5.69054pt\hskip-42.67912pt{\rm with}$ $\hskip 51.21504pt\xi\equiv\frac{|\dot{\phi}|}{2H\Lambda}\,.$ (21)
For values $\xi\gtrsim\mathcal{O}(1)$, one of the chiral modes develops an exponential growth, while the other remains in vacuum. For $\dot{\phi}<0$ (>0>0), it is $A^{+}$ ($A^{-}$) the one that experiences the instability. Considering $\xi$ to be approximately constant, as expected deep inside the slow-roll dynamics, a general solution for Eq. (20) is given by
@@
with $F_{0}$ and $G_{0}$ the regular and irregular Coulomb wave functions [11, 12]. The coefficients have been chosen so that for scales deep inside the Hubble radius, $k\tau\gg 1$, both chiralities are in the so-called Bunch-Davies (BD) vacuum,
@@
Without loss of generality, we choose $\dot{\phi}<0$, so that the excited mode is $\mathcal{A}^{+}$. In the limit $k\tau\ll 2\xi$, the solution (22) for this mode can be approximated by
@@
Depending on the evolution of $\xi$, the exponential growth of the unstable mode may become sufficiently large so that its contribution to the expansion of the universe (in the form of electromagnetic energy density) may become gradually more and more relevant. Furthermore, the term $\vec{E}\cdot\vec{B}\,$ in the rhsrhs of Eq. (5) may also become relevant, affecting the inflaton dynamics. In other words, for sufficiently large coupling $\alpha_{\Lambda}$, as inflation proceeds, there is always a point at which the backreaction-less approximation becomes unreliable. We must, therefore, refine the treatment of the dynamics, as we do next, in order to incorporate the backreaction of the gauge field.
@@ II.3 Local backreaction
$\hskip-19.91684ptH({\pi}_{\phi}^{\prime}+3\pi_{\phi})\hskip-2.84526pt$ $=$ $\hskip-2.84526pt\frac{1}{a^{2}}\nabla^{2}\phi-m^{2}\phi+\frac{\alpha_{\Lambda}}{a^{3}m_{p}}\vec{E}\cdot\vec{B}\,$, (31)
$\hskip-19.91684ptH\big(\vec{E}^{\prime}+\vec{E}\,\big)\hskip-2.84526pt$ $=$ $\hskip-2.84526pt\frac{1}{a^{2}}\vec{\nabla}\times\vec{B}-\frac{\alpha_{\Lambda}}{am_{p}}\big(\pi_{\phi}\vec{B}-\vec{\nabla}\phi\times\vec{E}\,\big),$ (32)
$\hskip-19.91684pt-3HH^{\prime}$ $=$ ${1\over m_{p}^{2}}(3\rho_{\rm K}+\rho_{\rm G}+2\rho_{\rm EM})\,$, (33)
@@ II.4 Chiral Projection
which, by construction, is antisymmetric, $\Sigma_{ij}=-\Sigma_{ji}$, and has $\vec{\varepsilon}^{\,\pm}$ as eigenvectors, with eigenvalues $\pm 1$, i.e. $\Sigma_{ij}\varepsilon_{j}^{\lambda}=\lambda\varepsilon_{i}^{\lambda}$. Extending the basis to a full orthonormal triad $\{\vec{\varepsilon}^{\,+},\vec{\varepsilon}^{\,-},\hat{k}\}$, and relabelling the vectors as $\vec{\varepsilon}^{\,(-1)}\equiv\vec{\varepsilon}^{\,-},\vec{\varepsilon}^{\,(0)}\equiv\hat{k},\vec{\varepsilon}^{\,(+1)}\equiv\vec{\varepsilon}^{\,+}$, we see that $\vec{\varepsilon}^{\,(0)}$ is also an eigenvector of $\Sigma_{ij}$ with vanishing eigenvalue, i.e. $\Sigma_{ij}\varepsilon_{j}^{\,(0)}=0$. We can therefore write
@@
[…]
$\hskip-28.45274ptA^{\pm}_{i}({\mathbf{k}})\equiv\Pi^{\pm}_{ij}(\hat{\mathbf{k}})A_{j}({\mathbf{k}})\,,~~E^{\pm}_{i}({\mathbf{k}})\equiv\Pi^{\pm}_{ij}(\hat{\mathbf{k}})E_{j}({\mathbf{k}})\,.$ (40)
@@ III Lattice formulation
We consider that $\phi$ lives on the lattice sites n, $A_{i}$’s on the links (i.e. halfway in between lattice sites, at $\textbf{n}+\hat{\imath}/2$), and we define spatial lattice derivatives as usual, using forward/backward operators $\Delta^{\pm}_{i}\varphi\equiv\pm\frac{1}{\delta x}(\varphi_{\pm i}-\varphi)$, where $\delta x$ is the lattice spacing, with the $\pm{i}$ subscripts indicating spatial displacements in one lattice unit in the direction $\hat{\imath}$. We define electric and magnetic fields on the lattice as
@@
[…]
@@
$\begin{gathered}\rho^{\rm L}_{\rm K}\equiv\frac{1}{2}\left\langle\pi^{2}_{\phi}\right\rangle_{\rm V}\;,\quad\rho^{\rm L}_{\rm G}\equiv\frac{1}{2a^{2}}\Big\langle(\vec{\Delta}^{\hskip-1.42262pt+}\hskip-1.42262pt\phi)^{2}\Big\rangle_{\rm V}\;,\\ \rho^{\rm L}_{\rm V}\equiv\frac{1}{2}m^{2}\left\langle\phi^{2}\right\rangle_{\rm V}\;,\quad\rho^{\rm L}_{\rm EM}\equiv\frac{1}{2a^{4}}\Big\langle a^{2}\vec{E}^{2}+\vec{B}^{2}\Big\rangle_{\rm V}\;,\end{gathered}$ (49)
[…]
@@ III.1 The reciprocal lattice
[…]
@@
[…]
The Fourier transform of these $A_{i}^{\lambda}(\tilde{\textbf{n}})\longrightarrow A_{i}^{\lambda}({\textbf{n}})$ will be transverse only with respect to the forward/background derivative operators, i.e. $\Delta_{i}^{\pm}A_{i}^{\lambda}({\bf n})=0$ [^1] Note that $\pm$ in $\Delta_{i}^{\pm}$ refers to forward (+) / backward (-) derivatives, unrelated to the chiral component $\lambda$, which are also either ++ or - (independently of the choice of derivative)..
@@
[…]
@@
$\langle{f}^{2}\rangle=\int\hskip-2.84526ptd\log k~\Delta_{\tt f}(k)\,,~~~\Delta_{f}(k)=\frac{k^{3}}{2\pi^{2}}\mathcal{P}_{f}(k)\,$,
@@
[…]
@@
with $_{l}$ the exact multiplicity of modes within the $l$-th bin, and $\langle\left|f(\tilde{\bf n}^{\prime})\right|^{2}\rangle_{R(l)}\equiv{1\over\#_{l}}\sum_{\tilde{\bf n}^{\prime}\in R(l)}\left|f(\tilde{\bf n}^{\prime})\right|^{2}$ an angular average over all modes $\tilde{\bf n}^{\prime}$ inside the $l$-th bin, represented by $R(l)\equiv\{\tilde{\bf n}^{\prime}~/~|\tilde{\bf n}^{\prime}|\in[l-1/2,l+1/2)\}$. We note that within the IR region ($l<N/2$) of the lattice, Eq. (60) is well approximated by $\Delta_{f}(k_{l})\simeq(k_{l}^{3}/2\pi^{2})(\delta x^{3}/N^{3})\large\langle|f(\tilde{\bf n}^{\prime})|^{2}\large\rangle_{R(l)}$, resembling closely the continuum form $(k^{3}/2\pi^{2})\mathcal{P}_{f}(k)$. Furthermore, the expectation value obtained with the lattice power spectrum just defined, $\left\langle f^{2}\right\rangle=\sum_{l=1}^{l_{\rm max}}{1\over l}\Delta_{f}(k_{l})$, is exactly identical to the lattice volume average $\left\langle f^{2}\right\rangle_{V}\equiv\frac{1}{N^{3}}\sum_{\bf n}f^{2}({\bf n})$, and mimics the continuum counterpart $\int d\log k~\Delta_{f}(k)$ as desired, since ${1/l}\equiv{k_{\rm IR}/k_{l}}\simeq dk/k\equiv d\log k$. In practice, we compute power spectra in our simulations for $\phi,\vec{A}^{\pm},\vec{A},\vec{E}^{\pm},\vec{E},\vec{B}$ as $\Delta_{\phi},\Delta_{A}^{(\pm)},\Delta_{A},\Delta_{E}^{(\pm)},\Delta_{E},\Delta_{B}$ via Eq. (60) with $|f(\tilde{\bf n})|^{2}=$ $|\phi(\tilde{\bf n})|^{2}$, $|\vec{A}^{\pm}(\tilde{\bf n})|^{2}$, $|\vec{A}(\tilde{\bf n})|^{2}$, $|\vec{E}^{\pm}(\tilde{\bf n})|^{2}$, $|\vec{E}(\tilde{\bf n})|^{2}$, and $|\vec{B}(\tilde{\bf n})|^{2}$, respectively, with the electric and magnetic fields given by Eq. (42).
[…]